I need help getting this problem solved. I need this by tomorrow at midnight. Please Help!
Select any imaginary number (of the form "a + bi," where a and b are non-zero real numbers), and another number such that the sum, difference, product, or quotient of the two numbers is a real number

Answer by stanbon(75887)

Select any imaginary number (of the form "a + bi," where a and b are non-zero real numbers), and another number such that the sum, difference, product, or quotient of the two numbers is a real number
--------------------------
Product: (2+3i)(2-3i) = 4 +6i - 6i -9i^2
= 4+9
= 13
======================
Cheers,
Stan H.